calculate the following nutrition needs for patty a 75 year old woman with long history of hypertension who is starting on peritoneal dialysis height 5’2” weight 150
Answer #1

Weight of the patient is 150lbs = 150×0.45kg= 67.5kg

Height of the patient is 5'2" = 157cm

Energy requirement :-

35 kcal/kg/day for patients younger than 65

30-35 kcal/kg/day for patients older than 65

= 30× 67.5 /day is the energy requirement

Protein requirement :-

1.2 g/kg/day

= 1.2× 67.5 = 81.2gm per day